At 382 C the reaction between astatine gas (diatomic) and hydrogen gas in equilibrium with monoastatine trihydrogen gas has a Kc = 85. What is the numerical value of Kp for this reaction? Please show all work including a balanced equation:
Answer #1

The balance equation is

We know

Kp=Kc*(RT)

Where

Kc=85

T =382 °C =273+382=655K

universal gas constant R =0.08206 L atm K-1mol-1

kp=85*(0.08206*655)-2=0.0307 ans